OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (anonymous):

Use the point slope form: Y-Y1=m(X-X1) m is your slope m=-1 and they gave you the points (1,-5). Then we just plug in the info in the point slope form: so Y-(-5)=-1(X-1). Y+5=-X+1, Then we solve for Y=-X-4. Hope I helped